OpenStudy (anonymous):

Find the slope of the following equation: 3x + 2y = 6 (5 points) Answer -3/2 2/3 3/2 -2/3 None of these

OpenStudy (mathhelp346):

By -3/2, etc., do you mean (-3,2)?

OpenStudy (aravindg):

rearrange in form y=mx+c

OpenStudy (aravindg):

where m is slope

OpenStudy (aravindg):

2y=6-3x y=-3/2x+3

OpenStudy (aravindg):

so slope=-3/2

OpenStudy (aravindg):

did that help?